OpenStudy (anonymous):

WILL FAN + MEDAL A diagonal of a parallelogram is 20 inches long and makes angles of 17° and 49° with the sides. How long is the longest side?

OpenStudy (tkhunny):

You will need a picture. Diagonals are perpendicular.

OpenStudy (anonymous):

6.4 in. 7.7 in. 16.5 in. 24.2 in.

OpenStudy (anonymous):

those are the choices

OpenStudy (tkhunny):

You'll still need a picture. You can use the [Draw] button below. So far, the question isn't quite making sense.

OpenStudy (anonymous):

It's 16.5 in.

OpenStudy (perl):



OpenStudy (perl):

you can use the fact that a triangle adds up to 180 17 + 49 + third_angle = 180 third_angle = 180 - 17-49 then you can use law of sines

OpenStudy (perl):



OpenStudy (perl):

$$ \Large { \frac{x}{\sin 17^o}=\frac{20}{\sin 114^o} \\~\\~\\ \frac{y}{\sin 49^o}=\frac{20}{\sin 114^o} } $$